i have a room 324 sq feet and i want to cover it with boards 1 x 8 x 10 how many boards will I need?
The room is 18 x 18 with 8 feet ceilings

Answer by Alan3354(69443)

i have a room 324 sq feet and i want to cover it with boards 1 x 8 x 10 how many boards will I need?
The room is 18 x 18 with 8 feet ceilings
--------------
I assume the boards are 8 inches by 10 feet, and 1 is the thickness ?? If that's not right, you'll have to supply dimensions with units.
----
Each board has an area of 8*10*12 = 960 sq inches.
324 sq ft = 324*144 sq inches = 46656 sq inches
46656/960 = 48.6 boards, IF you can cut them and put them together end to end without any loss (no sawdust).
-----
With some scrap, the number of boards needed will increase.